OpenStudy (anonymous):

what is the simplified form of the expression ? 5(14-2)^2 over 2 A.60 B.30 C.72 D.360

OpenStudy (here_to_help15):

may be ahead of me by 1 grade but i took this last year :P

OpenStudy (here_to_help15):

you

OpenStudy (anonymous):

okay i just need the answer haha

OpenStudy (here_to_help15):

k : )

OpenStudy (here_to_help15):

do u know PEMDAS?

OpenStudy (here_to_help15):

fine its 360

OpenStudy (here_to_help15):

but do u need help?

OpenStudy (igreen):

Is it this? \(\dfrac{5(14-2)^2}{2}\)

OpenStudy (here_to_help15):

yup : )

OpenStudy (igreen):

@Here_to_Help15 Its against the site's rules to give out direct answers without any explanation.

OpenStudy (here_to_help15):

lol i was nvm

OpenStudy (anonymous):

yes thats it

OpenStudy (here_to_help15):

LIKE I SAID I GREEN PEMDAS

OpenStudy (igreen):

\(\dfrac{5(14-2)^2}{2}\) PEMDAS, Parenthesis first: \(\dfrac{5(12)^2}{2}\) PEMDAS, Exponent next: \(\dfrac{5(144)}{2}\) PEMDAS, Multiplication next: \(\dfrac{720}{2}\) PEMDAS, lastly Division: \(360\)

OpenStudy (here_to_help15):

What did i say PEMDAS

OpenStudy (here_to_help15):

and so u know i was gonna show her how to get the answer until u rudely interupted i think that is probaly against the rules

OpenStudy (here_to_help15):

as well

OpenStudy (conqueror):

That is incorrect. @Here_to_Help15 Don't just provide the answer to a problem when someone else is in the middle of helping! But if you want to help, by all means, join in!

OpenStudy (conqueror):

That sentence is from the CoC.

OpenStudy (here_to_help15):

wtf i didnt i was the one here until he poped up

OpenStudy (here_to_help15):

read the damn story lol o whatever this thing is called

OpenStudy (here_to_help15):

youll see me asking her if she needed help

OpenStudy (igreen):

@Here_to_Help15 You're supposed to EXPLAIN FIRST, before you post your answer.

OpenStudy (here_to_help15):

well what if she didnt want the explanation

OpenStudy (igreen):

It's their rules.

OpenStudy (here_to_help15):

as you see she stated "okay i just need the answer haha" wtf does that tell you

OpenStudy (conqueror):

Well, I meant to post the last part of that sentence.

OpenStudy (here_to_help15):

i dont see anywhere in that sentence her asking for a explanation

OpenStudy (conqueror):

It doesn't matter if she doesn't want an explanation, she doesn't make the decision.

OpenStudy (igreen):

^ She will still need it no matter what.

OpenStudy (anonymous):

OK THANKS FOR THE HELP GUYS STOP ARGUING PLS THANKS OMG

OpenStudy (here_to_help15):

she does have the decision

OpenStudy (conqueror):

@lolpizza. Have homework problems? We can help you - as long as you are looking to learn how to get your solution (and not just the answer).

OpenStudy (here_to_help15):

no 1 else is controlling her or u or me with damn rules

OpenStudy (here_to_help15):

just messenge me when u need the answer

OpenStudy (conqueror):

Don't be offensive, inappropriate, or creepy! So avoid: Cursing and bypassing the language filters.

OpenStudy (here_to_help15):

or if you want the explanation go to mr igreen here

OpenStudy (here_to_help15):

i was glad to help lolpizza

OpenStudy (conqueror):

@shrutipande9

OpenStudy (igreen):

She will need an explanation from anybody who gives her an answer. The rules apply to everyone, and if you don't follow them you will be banned.

OpenStudy (conqueror):

It's pointless arguing with cheaters.

OpenStudy (here_to_help15):

lol theres always cheaters in life so get use to it